The calligra package^{∗} for use with L^{A}TEX2e

Gerd Neugebauer Mainzer Str. 16 56321 Rhens

             gerd@informatik.uni-koblenz.de

1996/07/18

1 Introduction

The font calligra provides a calligraphic font. Calligra can be found on the CTAN in the directory tex-archive/fonts/calligra. This package provides means to use this font in L^{A}TEX 2ε.

2 Usage

This file can be used as a package by placing its name in the argument of \usepackage. Afterwards the font family calligra is defined. This could also have been done by providing a font definition file.

The font definitions in this file scale down the original fonts to allow L^{A}TEX to choose the right baselineskip.

\calligra The command \calligra changes the current font family to calligra and the encoding to T1. Usually this should be used in a TEX group only since the macros \ss is redefined as well.

\textcalligra The command \calligra typesets its argument with the calligra font.

3 Implementation

First we declare a new font family for the calligra font.

 1 \DeclareFontFamily{T1}{calligra}{}

This font is only available in the normal shape. Here we can get the desired font by (silently) scaling the only present calligra15. The appropriate factor seems to be magstep 2 = 1.44.

 2 \DeclareFontShape{T1}{calligra}{m}{n}{<->s*[1.44]callig15}{}

Now we define the font changing command.

∗This file has version number v1.8, dated 1996/07/18.

\calligra The macro \calligra selects the calligra family. Since the kerning of the ß is terribly wrong we correct it within the macro \ss.

3 \DeclareRobustCommand\calligra{%
4 \renewcommand\ss{\symbol{255}\kern-.22em}%
5 \fontfamily{calligra}%
6 \fontencoding{T1}%
7 \selectfont}

\textcalligra The macro \textcalligra typesets its argument with the calligra font.

8 \newcommand\textcalligra[1]{\begingroup
9 \renewcommand\ss{\symbol{255}\kern-.22em}%
10 \fontfamily{calligra}%
11 \fontencoding{T1}%
12 \selectfont #1\endgroup}